OpenStudy (anonymous):

(x + 3) (x^2 - 4x + 16) can someone help me factor this with the sums of cubes pleaseeeeeeeeeeeee

OpenStudy (anonymous):

(x+3)(x^2-4x+16) =x^3-4x^2+16x+3x^2-12x+48 =x^3-x^2+14x+48